Suppose that the economy is characterised by the following behavioural equations:
C = C0 + c1Yd
Where:
C0 = 280
12
MAE206D /April 2020
I = 270
G = 300
T = 200
Marginal Propensity to Save (MPS) is 0.4 (or 40%)
Solve for:
a. equilibrium GDP (Y).
b. disposable income (Yd).
c. consumption spending (C).
1
Expert's answer
2020-04-27T07:45:20-0400

a. Y=C+I+G

Y=C0+C1Yd+I+G

Y=Co+C1(Y-T)+I+G

MPS+MPC=1,MPC=1-MPS=1-0.4=0.6

Y=280+0.6(Y-200)+270+300

Y=0.6Y+730

Y-0.6Y=730

0.4Y=730

Y=1825

b. Yd=Y-T=1825-200=1625

c. C=C0 + c1Yd=280+0.6(1825-200)=1255